### v2

# Gluon Condensate via Dirac Spectral Density:

Abstract
Quark and gluon scalar densities, $\langle\bar{\psi}\psi\rangle$ and $\langle F^{2}\rangle$, reflect the degree of scale-invariance violations in SU(N) gauge theories with fundamental quarks. It is known that $\langle\bar{\psi}\psi\rangle$ can be usefully scale-decomposed via spectral density $\rho(\lambda)$ of Dirac modes. Here I give such formula for $\langle F^{2}\rangle$, which reveals that gluon condensate is a strictly UV quantity. For the recently-found IR phase [1, 2], where the infrared (IR) degrees of freedom separate out and become independent of the system’s bulk, it implies that $\langle F^{2}\rangle$ due to this IR part vanishes. Its glue thus doesn’t contribute to scale anomaly of the entire system and is, in this sense, scale invariant consistently with the original claim. Associated formulas are used us to define IR decoupling of glue, which may serve as an alternative indicator of IR phase transition. Using the simplest form of coherent lattice QCD, we express the effective action of full QCD entirely via Dirac spectral density.
Keywords: IR phase, gluon condensate, QCD phase transition, quark-gluon plasma, scale invariance, scale anomaly, spectral density, IR-Bulk decoupling, coherent lattice QCD footnotetext: aihorv2@g.uky.edu
1. Introduction. Quark and gluon scalar densities, namely the expectation values $\langle\bar{\psi}\psi\rangle$ and $\langle F^{2}\rangle$, play an important role in the dynamics of QCD. For example, in hadronic physics they are well-known as leading “condensates” entering the QCD sum rules [3]. At a more basic level, they are connected to fundamental symmetries which makes them objects of prime interest for understanding the phases of nature’s strong interactions and, more generally, the phases in a large class of theories where quarks and gluons interact in the same manner. For purposes of the present discussion, this class $\cal T$} will consist of asymptotically-free vectorlike SU(3) gauge theories with fundamental quarks. Thus, $\cal T$} involves theories with arbitrary number $N_{f}\!<\!16.5$ of fundamental quark flavors with arbitrary masses $m_{q}$, $q=1,2,\ldots,N_{f}$ and at any temperature, including T=0T\!=\!0.
Viewed through the lens of scalar densities, each theory from $\cal T$} is characterized by $N_{f}\!+\!1$ values $\langle F^{2}\rangle$, $\langle\bar{\psi}\psi_{q}\rangle$. But even their hypothetical full knowledge wouldn’t provide us with understanding and classification of phases in the entire $\cal T$}. Indeed, the conventional symmetry-based apparatus for such analyses simply doesn’t have a full reach within $\cal T$}. One useful but limited classification is usually done within the subclass of theories containing multiple massless flavors, where $\langle\bar{\psi}\psi\rangle$ indicates the spontaneous breakdown of the associated flavored chiral symmetry. Similar goes for considerations concerning the anomalous nature of flavor-singlet chiral symmetry. While these special circumstances and their implications are important for understanding certain aspects of low-energy “real-world” QCD, they are in themselves unlikely to reveal the phase structure of the entire $\cal T$}. The situation is similar (but also different; see below) in case of dilations where full scale invariance requires quarks to be massless and temperature to be zero.
Figure 1: Types of thermal states for theories in $\cal T$} and schematics of their Dirac spectral densities $\rho(\lambda)$. Here $\lambda$ is the Dirac eigenvalue (scale) in the continuum-like notation where $D\psi_{\lambda}=i\lambda\psi_{\lambda}$. Left: B phase (standard confined phase) involves a single-component system with correlated parts. Its leading IR power behavior $\lambda^{0}$ includes cases when density is logarithmically divergent. Middle: IR phase involves a multi-component system with IR separated and decoupled from the bulk. $\Lambda_{\scriptscriptstyle{\text{\rm IR}}}$ is the energy scale of IR-bulk separation and $\lambda_{dc}$ the associated Dirac scale. Right: the hypothetical UV phase describes a single-component system of weakly interacting quarks and gluons.
A dramatic departure from the traditional and purely symmetry-based considerations appeared in Refs. [1, 2] which, together with refinements in Refs. [4, 5] and an additional state-of-the-art numerical evidence from Refs. [6, 7, 8], led to the classification of phases in the entire $\cal T$}. This primarily stemmed from the finding that systems in $\cal T$} can partition their degrees of freedom and become multicomponent. More precisely, there is a dynamical regime within $\cal T$} where deep infrared (IR) field fluctuations proliferate, separate out and decouple from the rest of finite-energy fundamental fields (bulk). Such IR-bulk separation, or property of being multi-component in general, doesn’t depend on symmetries and can occur anywhere in $\cal T$}. The new regime became known as the IR phase of $\cal T$} and was associated with restored scale invariance of glue in the IR component [2].
IR-bulk separation occurs in the Dirac space of theories in $\cal T$}, as shown schematically by Fig. 1 in the setting of thermal transitions [2]. The theory that “confines” at T=0T\!=\!0 (left), such as pure-glue QCD or real-world QCD, describes a single-component system at low $T$ (blue bulk) and features field fluctuations correlated over any pair of Dirac ($\lambda$) and physical energy ($E$) scales. At certain temperature $T_{\scriptscriptstyle{\text{\rm IR}}}$ thermal fluctuations become strong enough to induce the creation of independent IR component (red medium) whose degrees of freedom do not participate in the makeup of hadrons (“partial deconfinement”) and the IR phase ensues. Yet stronger disorder at $T\!>\!T_{\scriptscriptstyle{\text{\rm UV}}}\!>\!T_{\scriptscriptstyle{\text{\rm IR}}}$ may inhibit deep-IR ($\lambda\!\neq\!0$) degrees of freedom, with the system becoming single-component again, possibly as a “deconfined” perturbative bulk. It is worth noting that, unlike the first two cases, this classic scenario for weakly-coupled quark-gluon plasma has not yet been clearly identified in QCD and its existence is uncertain. Schematics of Dirac spectral density $\rho(\lambda)$ in Fig. 1 conveys how quark degrees of freedom are distributed across scales in the three regimes.
Since the separation of components is scale-based it is desirable, if at all possible, to think of scalar densities as composed of scale-dependent parts. Note that spectral density already partitions degrees of freedom by scale. This aspect was thus innately present in the arguments of original works on IR phase [1, 2], including arguments for scale invariance of the IR component. However, more formal treatment does need a scale decomposition of scalar densities since departures from scale invariance due to the glue field and the quark field of mass $m$ are quantified by (scale anomaly) [9, 10, 11]
$T_{\mu\mu}\,=\,\frac{\beta(g)}{2g}\,\langle F^{2}\rangle\,+\,\bigl{(}1+\gamma_{m}(g)\bigr{)}\,m\langle\bar{\psi}\psi\rangle$ (1)
and only scale decompositions of $\langle F^{2}\rangle$ and $m\langle\bar{\psi}\psi\rangle$ can isolate the contribution of a component. [^1] Notation here is standard, with $T$ the energy-momentum tensor and $\beta$, $\gamma_{m}$ the conventional RG functions. To that end, we point out here the usefulness of the following expressions
$m\langle{\bar{\psi}}\psi\rangle=m\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\;\frac{1}{\lambda+m}\;\,\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda)\qquad\qquad\;\langle\,F^{2}\,\rangle=\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\,\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\;\lambda\;\,\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda)\quad$ (2)
with $\lambda\!=\!\lambda_{\scriptscriptstyle{\mathrm{R}}}\!+\!i\lambda_{\scriptscriptstyle{\mathrm{I}}}$ a complex eigenvalue of lattice Dirac operator $D$ defining quark dynamics in a regularized theory (UV cutoff $1/a$; IR cutoff $1/L$). This $D$ is chosen such that the associated dimensionless constant $c_{\scriptscriptstyle{\mathrm{S}}}$ is non-zero (see below). Quantity $\rho_{s}(\lambda)\!=\!\langle n(\lambda,d{\cal S})\rangle/(V_{4}\,d{\cal S})$ is the surface spectral density of $D$, with $n(\lambda,d{\cal S})$ the number of eigenvalues in the area $d{\cal S}\!=\!d\lambda_{\scriptscriptstyle{\mathrm{R}}}d\lambda_{\scriptscriptstyle{\mathrm{I}}}$ around $\lambda$, and $V_{4}=L^{3}/T$. [^2] Notation ${\mathbb{R}}^{2}[{\mathbb{C}}]$ conveys that components in the domain ${\mathbb{R}}^{2}$ have meaning in the complex plane $\mathbb{C}$}. Dimension of $\rho_{s}(\lambda)$ is $a^{-2}$ instead of $a^{-3}$ in the standard definition. Effective density $\rho_{s}^{\scriptscriptstyle{\text{ef}}}\equiv\rho_{s}\!-\!\rho_{s0}$ subtracts that of the free field. [^3] We will see that in the glue case the free-field subtraction is necessary to merely define the operator. In the quark case it removes the leading UV divergence at non-zero $m$, which keeps the two expressions symmetric in this sense. Note that $\rho_{s}^{\scriptscriptstyle{\text{ef}}}=\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda,a,L)$ and $\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\;\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda)=0$.
Albeit not in this form, the quark relation in (2) appeared in considerations leading to Banks-Casher relation [12]. The glue expression is new and exploits the ideas put forward in Refs. [13, 14, 15]. Together they expose, in an explicit way, the key difference between the two scalar densities in terms of their scale makup: while the kinematic factor $1/(\lambda+m)$ enhances IR and suppresses UV in quark case, its counterpart $\lambda$ does the exact opposite in glue case. In fact, the glue expression invokes a surprising possibility that gluon condensate is an entirely UV quantity. We will show how Eqs. (2) clarify the proposed relation of IR phase to scale invariance [1, 2], and how to use the associated formulas to characterize its IR-bulk decoupling. We make all these results solid by writing the corresponding expressions (Eqs. (7) and (8)) for the important case of the overlap Dirac operator.
2. Gluon Condensate. We now derive the expression for gluon condensate in (2). To that end, consider the Euclidean lattice setup for theories in $\cal T$}, consisting of $N_{s}^{3}\times N_{\tau}$ sites of a hypercubic lattice with UV cutoff $1/a$, IR cutoff $1/L$ ($L\!=\!N_{s}a$), and temperature $T$ ($1/T\!=\!N_{\tau}a$). Lattice spacing $a$ is controlled by the gauge coupling $g$ but this relationship can be kept implicit here. Let $U\!\equiv\!\{U_{\mu}(x)\}$ be a configuration of SU(3) gauge field and D=D(U)D\!=\!D(U) a lattice Dirac operator. The latter is explicitly assumed to be local, gauge covariant and to respect hypercubic symmetries. Local gauge operators of definite space-time transformation properties can then be constructed from $D$ via suitable tracing operations [13, 14]. In case of $F^{2}(x)\equiv{\mathop{\rm tr}}_{c}\,F_{\mu\nu}F_{\mu\nu}(x)$ one starts from [13, 14, 15]
${\mathop{\rm tr}}_{cs}\hat{D}_{x,x}(U)-{\mathop{\rm tr}}_{cs}\hat{D}_{x,x}({\mathbb{I}})\,=\,c_{\scriptscriptstyle{\mathrm{S}}}\,a^{4}\,{\mathop{\rm tr}}_{c}\,F_{\mu\nu}F_{\mu\nu}(x,A)+{\cal O}(a^{6})\quad$ (3)
for transcription $U$ of a classical continuum gauge field $A$ onto the hypercubic lattice with spacing $a$, and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}+i\,[A_{\mu},A_{\nu}]$. Here $\mathbb{I}$} denotes the free field $U_{\mu}(x)\!\equiv\!\mathop{\rm diag}\{1,1,1\}$, ${\mathop{\rm tr}}_{c}$ is the trace in color while ${\mathop{\rm tr}}_{cs}$ in color-spin, and $c_{\scriptscriptstyle{\mathrm{S}}}$ is a constant. Note that $\hat{D}=aD$ is dimensionless but $F_{\mu\nu}$ has its physical dimension. [ In Refs. [13, 14, 15] $D$ denotes operator in lattice units.] Eq. (3) implies that, if $c_{\scriptscriptstyle{\mathrm{S}}}\!\neq\!0$, quantum operator $F^{2}(x)$ can be defined via indicated Dirac matrix elements, namely
$F^{2}(x,U)\equiv\frac{1}{c_{\scriptscriptstyle{\mathrm{S}}}\,a^{3}}\,{\mathop{\rm tr}}_{cs}\Bigl{[}D_{x,x}(U)-D_{x,x}({\mathbb{I}})\Bigr{]}\quad\,\longrightarrow\quad\,\,\langle\,F^{2}\,\rangle=\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\,\frac{T}{L^{3}}\,\Bigl{\langle}\,{\mathop{\rm Tr}}\Bigl{[}D(U)-D({\mathbb{I}})\Bigr{]}\,\Bigr{\rangle}\;$ (4)
where $\mathop{\rm Tr}$}” denotes the full trace of Dirac matrix. When performing the QCD average, we replaced $F^{2}(x)$ with $(\sum_{y}F^{2}(y))/(N_{s}^{3}N_{\tau})$, permitted by virtue of hypercubic translation invariance. For general Ginsparg-Wilson operators [16], non-zero $c_{\scriptscriptstyle{\mathrm{S}}}$ is expected due to their non-ultralocality [17, 18, 19]. In case of overlap Dirac operators [20] this was shown via explicit computation in Ref. [15].
Unlike in the continuum, eigenvalues $\lambda_{j}$ of lattice $D$ are not purely imaginary and details of the spectrum vary in different formulations. To cover all possibilities, we introduced in Sec. 1 the surface spectral density $\rho_{s}(\lambda)$ of eigenvalues in complex plane. This density is usefully represented as $\rho_{s}(\lambda,U)\!=\!\sum_{j}\delta(\lambda_{\scriptscriptstyle{\mathrm{R}}}-\lambda_{\scriptscriptstyle{\mathrm{R}}}^{j})\,\delta(\lambda_{\scriptscriptstyle{\mathrm{I}}}-\lambda_{\scriptscriptstyle{\mathrm{I}}}^{j})/V_{4}$ with $\lambda\!=\!\lambda_{\scriptscriptstyle{\mathrm{R}}}\!+\!i\lambda_{\scriptscriptstyle{\mathrm{I}}}$, $\lambda^{j}=\lambda^{j}(U)$ and $V_{4}=L^{3}/T$. Then
$\langle\,{\mathop{\rm Tr}}\,D\,\rangle\,=\,V_{4}\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\,\lambda\;\rho_{s}(\lambda)\qquad,\qquad\rho_{s}(\lambda)\equiv\langle\,\rho_{s}(\lambda,U)\,\rangle$ (5)
which consequently turns the expression for gluon condensate in Eq. (4) into one in Eq. (2) as desired.
3. Gluon Condensate with Overlap Operators. To analyze the implications of general Eqs. (2), it is desirable to examine them for operators $D$ that mimic continuum features to the largest extent possible. To that end, first note that neither quark nor gluon lattice condensate is guaranteed to be real-valued a priori. A natural way to ensure this is to impose $\gamma_{5}$-Hermiticity, namely $\gamma_{5}D\gamma_{5}=D^{\dagger}$, which forces eigenvalues to appear in complex-conjugate pairs, and physically amounts to continuum-like treatment of antiparticles. We will thus assume that $D$ is $\gamma_{5}$-Hermitian from now on.
It is also desirable to ensure continuum-like chiral properties on the lattice. At present, this amounts to using Ginsparg-Wilson (GW) Dirac operators [16]. [^4] See e.g. Refs. [18, 19] for full specification of the class. Among them, the 1-parameter family of overlap operators $D$ based on Wilson-Dirac matrix [20] has been studied most extensively. They are given by
$\frac{a}{\Delta}\,D\,=\,1+\frac{\hat{D}_{W}-\Delta}{\sqrt{\bigl{(}\hat{D}_{W}-\Delta\bigr{)}^{\dagger}\bigl{(}\hat{D}_{W}-\Delta\bigr{)}}}\quad\quad,\quad\quad\Delta\in(0,2)$ (6)
where $\hat{D}_{W}$ is the dimensionless massless Wilson-Dirac operator. In Ref. [15] it was shown that $c_{\scriptscriptstyle{\mathrm{S}}}\!=\!c_{\scriptscriptstyle{\mathrm{S}}}(\Delta)\!\neq\!0$ for the above $D\!=\!D(\Delta)$. They can thus be used to express gluon condensate via Eq. (2).
To that end, note that the $\gamma_{5}$-Hermitian spectrum of $D$ traces the circle of radius $\Delta$ centered at $(\Delta,0)$ so that the eigenvalues $\lambda_{\scriptscriptstyle{\mathrm{R}}}+i\lambda_{\scriptscriptstyle{\mathrm{I}}}$ satisfy $a(\lambda_{\scriptscriptstyle{\mathrm{R}}}^{2}+\lambda_{\scriptscriptstyle{\mathrm{I}}}^{2})=2\lambda_{\scriptscriptstyle{\mathrm{R}}}\Delta$. The expression involving spectral density along the circle is obtained via substituting $\rho_{s}(\sigma\cos\varphi,\sigma\sin\varphi)\!=\!\rho(\sigma)\delta(\varphi-\cos^{-1}(a\sigma/2\Delta))/\sigma$ upon transfer to polar coordinates in Eq. (2). We obtain
$\langle\,F^{2}\,\rangle_{a,L}\;=\;\frac{a^{2}}{c_{\scriptscriptstyle{\mathrm{S}}}\Delta}\,\int_{0}^{\bigl{(}\frac{2\Delta}{a}\bigr{)}^{-}}\!d\sigma\,\sigma^{2}\,\rho^{\scriptscriptstyle{\text{ef}}}(\sigma,a,L)\;\,+\;\,\,T\,\frac{\langle n_{0}\rangle_{a,L}}{L^{3}}\,\frac{2\Delta}{c_{\scriptscriptstyle{\mathrm{S}}}}\qquad$ (7)
where $\rho^{\scriptscriptstyle{\text{ef}}}\!=\!\rho-\rho_{0}$ has been defined previously and $\langle n_{0}\rangle$ is the average number of exact zeromodes equal to the number of modes with real eigenvalue $2\Delta/a$. While the former do not contribute to $\langle F^{2}\rangle$, the discrete contribution of the latter was separated out. In the continuum limit, energy-like variable $\sigma$ (magnitude of lattice eigenvalue) coincides with $\lambda$ of the continuum Euclidean formulation where eigenvalues are parametrized by $i\lambda$. Thus, $\rho(\sigma)$ is associated with the upper branch ($\lambda\!\geq\!0$) of the continuum density. Analogous derivation for the quark expression in (2) leads to
$m\langle{\bar{\psi}}\psi\rangle_{a,L}=\int_{0^{+}}^{\bigl{(}\frac{2\Delta}{a}\bigr{)}^{-}}\!d\sigma\;\frac{m^{2}+m^{2}+a\sigma^{2}m/\Delta}{m^{2}+\sigma^{2}+a\sigma^{2}m/\Delta}\;\,\rho^{\scriptscriptstyle{\text{ef}}}(\sigma,a,L)\;\,+\;\,\,T\,\frac{\langle n_{0}\rangle_{a,L}}{L^{3}}\,\Bigl{(}1+\frac{am}{2\Delta+am}\Bigr{)}\quad$ (8)
Here all real modes contribute and are separated out in the second term of the expression. We emphasize that Eqs. (7) and (8) are fully regularized scale decompositions of these quantities.
4. The Uses: Gluon Condensate as a UV Quantity. The above implies that, in a well-defined sense, gluon condensate in QCD is a UV quantity. The underlying logic is that the contribution of Dirac scales up to any finite renormalized value $\lambda_{\scriptscriptstyle{\mathrm{R}}}$ vanishes in the continuum limit. Indeed, writing the associated integral in Eq. (7) via renormalized quantities $\sigma_{\scriptscriptstyle{\mathrm{R}}}=\sigma/Z_{\scriptscriptstyle{\mathrm{S}}}$, $m_{\scriptscriptstyle{\mathrm{R}}}=m/Z_{\scriptscriptstyle{\mathrm{S}}}$, $\rho_{\scriptscriptstyle{\mathrm{R}}}(\sigma_{\scriptscriptstyle{\mathrm{R}}})\!=\!Z_{\scriptscriptstyle{\mathrm{S}}}\,\rho(Z_{\scriptscriptstyle{\mathrm{S}}}\sigma_{\scriptscriptstyle{\mathrm{R}}})$ [21, 22] we obtain
$\langle\,F^{2}\,\rangle_{L}\bigl{[}\lambda_{\scriptscriptstyle{\mathrm{R}}}\bigr{]}\;=\;\lim_{a\to 0}\,\frac{a^{2}Z_{S}^{2}}{c_{\scriptscriptstyle{\mathrm{S}}}\Delta}\,\int_{0}^{\lambda_{\scriptscriptstyle{\mathrm{R}}}}\!d\sigma_{\scriptscriptstyle{\mathrm{R}}}\,\sigma_{\scriptscriptstyle{\mathrm{R}}}^{2}\,\rho_{\scriptscriptstyle{\mathrm{R}}}^{\scriptscriptstyle{\text{ef}}}(\sigma_{\scriptscriptstyle{\mathrm{R}}},a,L)\;=\;0\quad\;,\quad\;\forall\;\,0<\lambda_{\scriptscriptstyle{\mathrm{R}}}<\infty\quad$ (9)
where the notation $\langle\,F^{2}\,\rangle[\lambda_{\scriptscriptstyle{\mathrm{R}}}]$ means discarding all bare Dirac scales whose renormalized counterpart exceeds $\lambda_{\scriptscriptstyle{\mathrm{R}}}$. The limit is zero since $\rho_{\scriptscriptstyle{\mathrm{R}}}^{\scriptscriptstyle{\text{ef}}}$ as well as the associated integral have a well-defined continuum limit at fixed $L$, and since the $a$-dependence of $Z_{\scriptscriptstyle{\mathrm{S}}}$ is at most logarithmic. Note also that the real modes, separated in Eq. (7), contribute at the scale of UV cutoff and thus do not enter Eq. (9).
Two additional remarks are worth making. Firstly, the above argument was made in the context of a finite system of size $L$. Regarding the infinite system, the same reasoning and conclusion applies for the order of limits $\lim_{L\to\infty}\lim_{a\to 0}$. For the reversed order, the first limit $\lim_{L\to\infty}$ is finite but $a$-dependent, and could in principle lead to such an UV divergence of the integral in (9) that would produce a non-zero contribution to gluon condensate. However, such UV divergence is unlikely to occur since it would have to be generated by IR Dirac modes. We thus propose that vanishing of any IR contribution to gluon condensate holds for both ways of removing cutoffs.
Secondly, note that the situation is very different for the quark part of trace anomaly. At finite fixed $L$ we have from Eq. (8) upon taking the continuum limit
$m\langle{\bar{\psi}}\psi\rangle_{L}\bigl{[}\lambda_{\scriptscriptstyle{\mathrm{R}}}\bigr{]}\,=\,2\int_{0^{+}}^{\lambda_{\scriptscriptstyle{\mathrm{R}}}}\!d\sigma_{\scriptscriptstyle{\mathrm{R}}}\,\frac{1}{1+\sigma_{\scriptscriptstyle{\mathrm{R}}}^{2}/m_{\scriptscriptstyle{\mathrm{R}}}^{2}}\;\rho^{\scriptscriptstyle{\text{ef}}}(\sigma_{\scriptscriptstyle{\mathrm{R}}},L)\;\,+\;\,T\,\frac{\langle n_{0}\rangle_{L}}{L^{3}}\quad$ (10)
where the second term is due to zero modes. The IR contribution is clearly non-zero with the second term vanishing in the $L\to\infty$ limit. The above form also shows that the quark contribution to trace anomaly becomes a strictly IR quantity in the chiral limit, which is familiar from considerations leading to Banks-Casher relation [12].
Figure 2: Left: schematic view of phases in set $\cal T$} based on a degree of deep-IR degrees of freedom proliferation and their scale invariance. Direction of arrows for parameters indicates the direction of possible phase changes along the chain $\text{B}\!\to\!\text{IR}\!\to\!\text{UV}$ [2]. Right: the case of near-massless quarks with Banks-Zaks (B-Z) regime and the asymptotic freedom (AF) boundary indicated.
5. The Uses: IR Scale Invariance in IR Phase. In the original work [2], IR phase was defined by the negative power in IR behavior of Dirac spectral density, and thus by a power-law enhancement of deep-IR degrees of freedom. The associated classification of phases in set $\cal T$} is
$\rho(\lambda)\propto\lambda^{p}\;,\;\lambda\to 0\quad\;\,\Longrightarrow\quad\;\,\text{phase}\;\,=\;\,\text{B}\;\,\text{if}\;\;p=0\;\;\;,\;\;\;\text{IR}\;\,\text{if}\;\;p<0\;\;\;,\;\;\;\text{UV}\;\,\text{if}\;\;p>0\;\;\;$ (11)
where continuum-like notation ($\sigma\to\lambda$) was used but the classification is already well-defined on the lattice. In IR phase, $p$ is usefully written as $p\!=\!-1\!+\!\delta$ because, at least in thermal cases, $\delta\!>\!0$ is small and may vanish in the continuum limit [2]. The “B” in B phase refers to broken IR scale invariance of glue: the standard IR characteristic of “confined phase”. Here spectral density is expected to logarithmically diverge away from the chiral limit [23, 2], and thus p=0p\!=\!0. In not-yet-observed UV phase, IR degrees of freedom are power-law suppressed with the power possibly infinite if spectral density is zero in deep IR. The associated phase structure of $\cal T$} is shown in Fig. 2 (left).
Two aspects [2] of the new IR phase are especially relevant for the present discussion. [^5] See Ref. [24] for more detailed discussion of features that make p<0p\!<\!0 regime a truly distinct phase. (i) Due to p<0p\!<\!0 in IR and p=3p\!=\!3 in UV (asymptotic freedom), with an intervening regime of severe mode depletion, $\rho(\lambda)$ of theories in IR phase exhibits a bimodal structure with the IR part separated from the bulk in scale/energy. This invoked the conjecture of IR-bulk decoupling, namely that the IR part is independent from the bulk and acts as an autonomous subsystem. (ii) Due to the near-pure negative-power behavior of $\rho$ and the connection to conformal window [2, 25, 26] (see Fig. 2 (right)), it was conjectured that glue of the IR part is scale invariant, at least asymptotically. Thus, although defined via Dirac spectral properties, the physics behind the above structure of $\cal T$} was proposed to be driven by glue and interpreted in glue terms [1, 2]. In that vein, the new Eq. (7) is perhaps the purest expression of the implied connection and allows us to express the key properties of IR phase, such as (i) and (ii), in more standard field-theoretic terms.
To that end, it is important to realize that (i) and (ii) are in fact connected [2, 4]. For example, while the IR-bulk separation is very suggestive of IR-bulk decoupling, it is not sufficient. But (ii) offers a dynamical reason for it which is easy to see in thermal IR phase at $N_{f}\!=\!0$. Indeed, assume that, upon entering the IR phase, this theory of glue becomes scale invariant below some energy $\Lambda_{\scriptscriptstyle{\text{\rm IR}}}<T$. The gauge coupling then stops running at that scale, which requires non-analyticities in the internal structure of the theory. These non-analyticities can then also facilitate the IR-bulk decoupling. Conversely, non-analyticities generated by breaking-off the IR into a physically distinct component can make scale invariance in IR possible. [^6] Thermal onset of these internal non-analyticities is predicted to occur at $T\!=\!T_{\scriptscriptstyle{\text{\rm IR}}}$ of IR phase transition [4, 5] because they induce non-analytic $T$-dependence of physical observables at that point.
There are at least two aspects of scale invariance to study in these IR phase circumstances. The first one isolates the IR component as a field system, bringing $\Lambda_{\scriptscriptstyle{\text{\rm IR}}}\rightarrow\infty$ by virtue of an overall rescaling, and aims to study scale-invariant field theory so defined, e.g. in the strongly-coupled part of conformal window. The second one views the multi-component system as a whole and seeks to understand the role its IR part plays in violations of scale invariance. Focusing on this second aspect of (ii), let $\lambda_{dc}=\lambda_{dc}(a)$ be the scale marking the IR-bulk boundary in the Dirac spectrum (see Fig. 1). The contribution $\langle F^{2}\rangle_{\scriptscriptstyle{\text{\rm IR}}}$ of the IR component (IR medium) to full $\langle F^{2}\rangle$ is
$\langle\,F^{2}\,\rangle_{{\scriptscriptstyle{\text{\rm IR}}}}\;\equiv\;\,\frac{a^{2}}{c_{\scriptscriptstyle{\mathrm{S}}}\Delta}\,\int_{0}^{\lambda_{dc}(a)}\!d\sigma\,\sigma^{2}\,\rho^{\scriptscriptstyle{\text{ef}}}(\sigma,a)\quad\longrightarrow\quad\,0\quad\;\,\text{for}\quad\;\,a\;\longrightarrow\;0\quad$ (12)
Its approach to zero in the continuum limit ensues due to obvious integrability at any finite $a$, and because $\lambda_{dc}(a)$ varies at most logarithmically in the vicinity of a=0a\!=\!0. Hence, the IR component of the system in IR phase doesn’t contribute to scale anomaly and is, from this point of view, scale invariant. The above argument is generic for IR phase of theories in $\cal T$}, and doesn’t depend on whether $L$ is kept fixed (finite) or taken to infinity at each $a$.
6. The Uses: IR-Bulk Decoupling in IR Phase. Aspects of the present analysis, and those of Refs. [13, 14, 15], can also be used to study IR-bulk decoupling in IR phase (point (i) in Sec. 5). The chief idea is that “decoupling” is identified with “decorrelation” which turns such parts into independent subsystems. Here we analyze the glue part of the system which is our focus in this work. The full account of spectral correlations and decoupling will be given elsewhere.
Since the IR-bulk separation is based on Dirac scales, we need the notion of correlation among such scale-based parts. This requires analogues of expressions in previous sections but for a given gauge background $U$. To that end, let’s consider the action-like dimensionless construct ${\cal F}^{2}(U)$
${\cal F}^{2}(U)\,\equiv\,a^{4}\sum_{x}F^{2}(x,U)\,=\,\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\,{\mathop{\rm Tr}}\Bigl{[}D(U)-D({\mathbb{I}})\Bigr{]}\,=\,V_{4}\,\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\,\lambda\;\rho_{s}^{\scriptscriptstyle{\text{ef}}}(\lambda,U)\quad$ (13)
written here for general $D(U)$ with $c_{\scriptscriptstyle{\mathrm{S}}}\!\neq\!0$, and where Eqs. (4) and (5) and their notation were used. For overlap Dirac operators we then have a specific expression (analogous to (7)), which allows to quantify the contribution to ${\cal F}^{2}(U)$ from range of lattice Dirac scales $\sigma$ ($0\leq\eta_{1}\leq\sigma\leq\eta_{2}<2\Delta/a$) as
${\cal F}^{2}(U,\eta_{1},\eta_{2})\,\equiv\,\frac{V_{4}a^{2}}{c_{\scriptscriptstyle{\mathrm{S}}}\Delta}\,\int_{\eta_{1}}^{\eta_{2}}\!d\sigma\,\sigma^{2}\,\rho^{\scriptscriptstyle{\text{ef}}}(\sigma,U)\;=\;\frac{a^{2}}{c_{\scriptscriptstyle{\mathrm{S}}}\Delta}\,\sum_{\eta_{1}<\sigma_{j}\leq\eta_{2}}\Bigl{[}\,\sigma_{j}^{2}(U)-\sigma_{j}^{2}({\mathbb{I}})\Bigr{]}\quad$ (14)
where the filter only allows values $\sigma_{j}(U)$ and (separately) $\sigma_{j}({\mathbb{I}})$ in the specified range. Note that the formula for given non-negative $\eta_{1}$, $\eta_{2}$ in fact combines (equal) contributions from both upper and lower branch of the Dirac spectrum. It is fully regularized and allows for defining correlations amongst different “parts” of the system with “partitions” based on Dirac spectra.
To formulate IR-Bulk decoupling of glue in IR phase, it is practical to make two preparatory steps. First, it is convenient to replace the variable $\sigma$ with $t\!=\!\sigma/m$, where $m$ may be e.g. the smallest non-zero quark mass in the theory, because it doesn’t get renormalized [21, 22]. Secondly, rather than the usual connected correlation $\langle{\cal O}_{1}{\cal O}_{2}\rangle_{\text{c}}\!\equiv\!\langle{\cal O}_{1}{\cal O}_{2}\rangle-\langle{\cal O}_{1}\rangle\langle{\cal O}_{2}\rangle$, we consider the normalized connected Pearson correlation $\langle\ldots\rangle_{\text{nc}}$ (covariance divided by standard deviations), namely
$\langle{\cal O}_{1}{\cal O}_{2}\rangle_{\text{nc}}\!\equiv\!\frac{\langle{\cal O}_{1}{\cal O}_{2}\rangle_{\text{c}}}{\sqrt{\langle{\cal O}_{1}{\cal O}_{1}\rangle_{\text{c}}\langle{\cal O}_{2}{\cal O}_{2}\rangle_{\text{c}}}}$ (15)
which scales out magnitudes of correlated observables and thus expresses pure correlation.
Let $t_{\text{ref}}$ be an arbitrary but fixed Dirac scale inside the bulk of the target continuum theory in IR phase defined by Eq. (11). We note in passing that the boundary between IR and bulk in IR phase may be defined by the existence of a point $t_{3}\!=\!\sigma_{3}/m\!>\!0$ in Dirac spectrum of a UV-regularized theory such that the spatial IR dimension of Dirac eigenmodes at $t\!>\!t_{3}$ is the ordinary $d_{\scriptscriptstyle{\text{\rm IR}}}\!=\!3$, while $d_{\scriptscriptstyle{\text{\rm IR}}}\!=\!0$ for modes at $0\!<\!t\!<\!t_{3}$ [4, 5, 7]. With that, $t_{\text{ref}}>t_{3}(a)$ for all sufficiently small $a$. IR-Bulk decoupling in IR phase then may be formulated as follows. In IR phase there exists at least one $0\!<\!\tilde{t}\!<\!t_{\text{ref}}$ such that the corresponding partitions of ${\cal F}^{2}(0,t_{\text{ref}})$ decorrelate, and the decoupling scale $t_{\text{dc}}$ is the largest of such scales, namely
$\lim_{a\to 0}\lim_{L\to\infty}\;\langle\,{\cal F}^{2}(0,\tilde{t})\,{\cal F}^{2}(\tilde{t},t_{\text{ref}})\,\rangle_{\text{nc}}\,=\,0\quad\quad\text{and}\quad\quad t_{\text{dc}}\,\equiv\,\sup\{\,\tilde{t}\,\}\quad$ (16)
Note that segment $(\tilde{t},t_{\text{ref}})$ contains part of bulk for each $\tilde{t}$.
Few remarks are important here. (i) It is implicitly understood that the above definition doesn’t depend on the choice of $t_{\text{ref}}$. (ii) Regarding $t_{\text{dc}}$, the available evidence [4, 5, 6, 7] embodied in the metal-to-critical scenario [5] favors the simplest possibility that $t_{\text{dc}}\!=\!t_{3}\!=\!t_{A}$ (see also Ref. [24]), where $t_{A}\!>\!0$ is the Anderson-like point of Refs [27, 28, 29]. However, studies directly confirming the above form of IR-Bulk decoupling, as well as direct computation of $t_{\text{dc}}$ are yet to be performed. (iii) It is consistent with the current knowledge that decoupling Eq. (16) may occur already before taking the continuum limit, i.e. even at non-zero $a$. (iv) Detection of IR phase (or even its definition) via the presence of the above IR-Bulk decoupling may be a fruitful approach to its numerical investigation. (v) Normalized Pearson correlations of ${\cal F}^{2}$ segments, such as those entering Eq. (16) can be expressed via via spectral values $\sigma_{i}(U)$ alone. The general expression is
$\big{\langle}\,{\cal F}^{2}(\eta_{1},\eta_{2})\,{\cal F}^{2}(\gamma_{1},\gamma_{2})\,\big{\rangle}_{\text{nc}}\,=\,\Big{\langle}\,\sum_{\eta_{1}<\sigma_{i}\leq\eta_{2}}\,\sigma_{i}^{2}(U)\,\sum_{\gamma_{1}<\sigma_{j}\leq\gamma_{2}}\,\sigma_{j}^{2}(U)\,\,\Big{\rangle}_{\text{nc}}\quad$ (17)
as follows from Eqs. (14) and (15), or in the same way via spectral $t$-values.
7. The Uses: Coherent Lattice QCD. The term coherent lattice QCD refers to formulations of lattice-regularized QCD constructed from a single object, namely a suitable lattice Dirac operator. Such theories were first proposed in Refs. [13, 14] and their construction utilized the same ideas as those employed here. The formulas involving Dirac spectral density, suggested in this work, provide an additional insight in their construction that we now make explicit.
Let $D$ be any lattice Dirac operator with $c_{\scriptscriptstyle{\mathrm{S}}}\!\neq\!0$. The basic coherent lattice QCD has the form
$S\,=\,\frac{1}{2g^{2}}\,\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\,{\mathop{\rm Tr}}\Bigl{[}D(U)-D({\mathbb{I}})\Bigr{]}\,+\,a^{4}\sum_{q=1}^{N_{f}}{\bar{\psi}}_{q}\Bigl{(}D(U)+m_{q}\Bigr{)}\psi_{q}$ (18)
where $N_{f}$ is the number of quark flavors with masses $m_{q}$. In original Refs. [13, 14] this action was written in terms of dimensionless lattice objects and factors involving powers of $a$ were thus absent. Note that the gauge part is in fact ${\cal F}^{2}(U)/2g^{2}$ with ${\cal F}^{2}$ given in Eq. (13).
The free-field part of the glue action is a constant that can be omitted in the definition of the theory. The effective glue action after integrating out the quark variables reads
$\begin{split}S_{\text{eff}}(U)&\,=\,\frac{1}{2g^{2}}\,\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\,{\mathop{\rm Tr}}\,D(U)\,-\,\sum_{q=1}^{N_{f}}\,{\mathop{\rm Tr}}\,\ln\big{(}aD(U)+am_{q}\big{)}\,=\,\\ &\,=\,V_{4}\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\,\rho_{s}(\lambda,U)\,\Bigl{[}\,\frac{1}{2g^{2}}\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\,\lambda\,-\,\sum_{q=1}^{N_{f}}\ln(a\lambda+am_{q})\,\Bigr{]}\end{split}$ (19)
where $\lambda$ is a complex variable and the notation was introduced in connection with Eq. (2). Thus, the defining object of the theory, the action, is scale-decomposed and expressed in terms of Dirac spectral density. There is another hidden constant in the above expression, isolated by factoring $am_{q}$ in log terms. Discarding it puts the effective action into the form
$S_{\text{eff}}(U)\,=\,V_{4}\int_{{\mathbb{R}}^{2}[{\mathbb{C}}]}d{\cal S}\,\rho_{s}(\lambda,U)\,\Bigl{[}\,\frac{1}{2g^{2}}\frac{a}{c_{\scriptscriptstyle{\mathrm{S}}}}\,\lambda\,-\,\sum_{q=1}^{N_{f}}\ln\big{(}1+\lambda/m_{q}\big{)}\,\Bigr{]}\,\equiv\,S_{\text{eff}}^{\scriptscriptstyle{G}}(U)+S_{\text{eff}}^{\scriptscriptstyle{Q}}(U)$ (20)
For the family of overlap Dirac operators we then have in particular for glue and quark parts
$2g^{2}\,S_{\text{eff}}^{\scriptscriptstyle{G}}(U)\,=\,\frac{2\Delta}{c_{\scriptscriptstyle{\mathrm{S}}}}\,n_{0}(U)\,+\,\frac{a^{2}V_{4}}{c_{\scriptscriptstyle{\mathrm{S}}}\Delta}\int_{0}^{\bigl{(}\frac{2\Delta}{a}\bigr{)}^{-}}\!d\sigma\,\rho^{\scriptscriptstyle{\text{ef}}}(\sigma,U)\,\sigma^{2}$ (21)
$S_{\text{eff}}^{\scriptscriptstyle{Q}}(U)\,=\,n_{0}(U)\sum_{q=1}^{N_{f}}\ln\Bigl{(}1\,+\,\frac{2\Delta}{am_{q}}\Bigr{)}\,+\,V_{4}\,\int_{0^{+}}^{\bigl{(}\frac{2\Delta}{a}\bigr{)}^{-}}\!d\sigma\,\rho^{\scriptscriptstyle{\text{ef}}}(\sigma,U)\,\sum_{q=1}^{N_{f}}\,\ln\Biggl{[}1+\frac{\sigma^{2}}{m_{q}^{2}}\Bigl{(}1+am_{q}/\Delta\Bigr{)}\Biggr{]}$ (22)
where the contribution of real modes was again separated. Construction of other coherent lattice QCD actions, such as the symmetric logarithmic case [13, 14], proceeds along the same lines.
8. Epilogue. Dirac spectral density $\rho(\lambda)$ in QCD specifies the distribution of its quark degrees of freedom over Dirac scales. Its knowledge thus facilitates, among other things, our understanding of how these scales contribute to the composition of important quark observables such as $\langle\bar{\psi}\psi\rangle$. For example, we learned that $\langle\bar{\psi}\psi\rangle$ is a strictly IR quantity in the chiral limit, which is a particular way to interpret the approach of Banks and Casher [12].
Here we showed, via Eqs. (2) and (7), that $\rho(\lambda)$ also determines how scalar glue density gets aportioned across these scales. The new formulas reveal that glue density, in a stark contrast to quark density, is a strictly UV quantity. As such, it also provides for the analogue of Banks-Casher relation for gluon condensate. Details will be discussed in a dedicated account, but our reasoning makes it clear that this quantity is encoded in $1/\lambda$ power term in UV asymptotics of $\rho(\lambda)$.
The outlook on $\rho(\lambda)$ as the distribution of DOFs over scales was pervasive in works that led to the discovery of IR phase [1, 2, 4, 5]. Definition of the phase is also expressed via IR asymptotics of $\rho(\lambda)$ [2]. But its key features, such as scale invariance and IR-Bulk separation, were always chiefly attributed to gluonic rather than quark degrees of freedom. This was not contradictory with the explanation that matrix elements of the Dirac operator are in fact gauge-covariant glue operators, which makes $\rho(\lambda)$ the unusual (scale-dependent) but glue operator as well. Here we made the needed direct link between Dirac spectral density and the aforementioned driving glue effects.
Our formulas have several applications, particularly in relation to IR phase where these considerations first arose. Here we discussed the question of glue scale invariance of the IR part and the precise definition of IR-Bulk decoupling. They are all based on our formulas featuring Dirac spectral density, which also possibly opens new ways of numerical evaluation. Our considerations are most potent in regularizations known as coherent lattice QCD [13, 14], whose full effective actions can also be expressed via Dirac spectral density.
Acknowledgments. Discussions with Vladimír Balek and Robert Mendris on the topics of this work are gratefully acknowledged. Kudos to Hali for her gentle support. Many thanks to Dimitris Petrellis for help with technical aspects. I may never be able to repay my debts to Sylvia and Vlado. This work was supported in part by US Department of Energy under grant DE-FG02-95ER40907.